# Memory-Enhanced Neural Solvers for Efficient Adaptation in Combinatorial Optimization

### v2

@@ 5 Conclusion
We present MEMENTO, a method to improve adaptation of neural CO solvers to unseen instances. MEMENTO enables to condition a policy directly on data collected online during search, facilitating rapid adaptation and efficient use of the search budget. In practice, it proves to outperform stochastic sampling, tree search and policy-gradient fine-tuning of the policy, and shows it can combined with another solver than the one used during training, with no further re-training needed. Additionally, MEMENTO respects several key properties, like robustness to low-budget regimes, and favorable time performance scalability, both in instance size and base solver size.
We demonstrate the efficacy of MEMENTO by achieving state-of-the-art single-policy adaptation on a standard benchmark on TSP and CVRP, both in and out-of-distribution. Moreover, we find that MEMENTO realises interpretable update rules to the underlying policy; trading off exploration and exploitation over the search budget and outperforming REINFORCE-style updates. We further demonstrate the flexibility of this approach by achieving zero-shot combination of MEMENTO and COMPASS, achieving overall state-of-the-art on a benchmark of $12$ tasks. The simplicity and efficacy of this method makes it amenable to a very large set of problems, laying another stone towards the application of neural solvers to real-world problems.
Limitations and future work. MEMENTO incurs additional compute and memory-usage compared to the memory-less base policies which it augments. Practically, we find that the performance gain significantly outweighs these overheads. A potential mitigation could be to derive the mathematical update learned by MEMENTO to avoid relying on the MLP computation. The adaptation enabled by MEMENTO is determined by the data collected by the base policy, therefore this could limit the efficiency when evaluation instances are very far from the training distribution, or when the base policy has harmfully poor initial performance. Finally, we believe that the capacity of MEMENTO could be improved by training on broader distributions, in particular using variable instance sizes.
@@ Appendix A Extended results
In Section 4, we compare MEMENTO to popular methods from the literature in numerous settings. In particular, Figure 3 compares using stochastic sampling, MEMENTO and EAS to adapt POMO on the standard benchmark; and Figure 6 compares the relative performances of MEMENTO and EAS on larger instances of TSP and CVRP for several values of batch sizes and budget. In this section, we report additional values and results of other methods for those experiments.
@@ A.1 Standard benchmark
We compare to the heuristic solver LKH3 (Helsgaun, 2017); the current leading industrial solver of both TSP and CVRP, as well as an exact solver Concorde (Applegate et al., 2006) which is a TSP-specific industrial solver, and the CVRP-specific solver (Vidal et al., 2012).
We use datasets of 10,000 instances with 100 cities/customer nodes drawn from the training distribution, and three generalization datasets of 1,000 instances of sizes 125, 150, and 200, all from benchmark sets frequently used in the literature (Kool et al., 2019; Kwon et al., 2020; Hottung et al., 2022; Grinsztajn et al., 2023; Chalumeau et al., 2023b). Tables  2(a) and  2(b) display results for TSP and CVRP on the standard benchmark, respectively. The columns provide the absolute tour length, the optimality gap, and the total inference time that each method takes to solve one instance within the attempts budget.
@@
We were able to do so on the standard benchmark since we have implementations of most methods we were comparing in a single framework. Since we are reporting several external methods in Fig. 7(a), we could only report time taken for the full dataset in that case.
@@
On Table 2, we can see that in the single-agent setting, MEMENTO leads the entire benchmark, and in the population-based setting, MEMENTO (COMPASS) is leading the whole benchmark. MEMENTO is able to give significant improvement to COMPASS on CVRP, pushing significantly the state-of-the-art on this benchmark. On TSP, the improvement is not significant enough to be visible on the rounded results. To finish with, we can observe that the time cost associated with MEMENTO is reasonable and is worth the performance improvement (except maybe for TSP results of MEMENTO (COMPASS)).
@@
All neural methods reported in Table 2 are our own runs with standardised checkpoints and rollout strategies, except for the results of SGBS, which were taken from Choo et al. (2022). This introduces three biases: (i) the base POMO checkpoint used by SGBS is not exactly the same as our re-trained POMO checkpoint (ii) SGBS uses the domain-specific augmentation trick that we do not use (iii) SGBS pre-selects starting points in CVRP, which we do not do.
@@
Table 2: Results of MEMENTO against the baseline algorithms for (a) TSP, (b) CVRP. The methods are evaluated on instances from training distribution ($n=100$) as well as on larger instance sizes to test generalization. We use the same dataset as the rest of the literature, those contain 10 0001000010\,00010 000 instances for $n=100$ and $1000$ instances for $n=125,150,200$. Gaps are computed relative to LKH3. We report time needed to solve one instance. *SGBS results are reported from Choo et al. (2022), and do not use the same POMO checkpoint as other reported results. Additionally, they rely on problem-specific tricks that were not used by other methods. Details in Section A.1.
@@
460S
@@ A.2 Performance over different number of parallel and sequential attempts
In Section 4.3, we show performance improvement of MEMENTO over EAS on instances of size 500 on TSP and CVRP for increasing number of sequential attempts and size of attempt batch. We report extended results with an additional competitor, POMO. We compare the online adaptation of the methods over four different sizes of batched solutions across increasing sequential attempts for each CO problem. Results from Tables 3(a) and 3(b) show results of the three methods on TSP and CVRP, respectively. The columns show the best tour length performance for various values of sequential attempts expressed as budget, and solution batches of size $N$. Performance is averaged over a set of 128 instances.
@@ A.3 Evaluation over larger instances
In Section 4.3, we evaluate MEMENTO and baselines on instances of size 500. For TSP, we use the dataset from Fu et al. (2021). For CVRP, we use the dataset from Luo et al. (2023). We do not include LEHD in our results since we focus on methods trained with Reinforcement Learning, and LEHD can only be successfully trained with supervised learning at the time of writing. Nevertheless, the good performance achieved by LEHD has motivated us to study the scaling law of MEMENTO as the number of layers in the decoder increases, reported on Fig. 7.
@@
A.4 Time complexity analysis
To get the time curves reported in Section 4.3, we used CVRP, since it is the environment were MEMENTO was the slowest compared to EAS. We hence expect curves on TSP to be even better for MEMENTO. For the instance size scaling, we use 50 starting points and 100 sequential attempts, and 1 layer in the decoder, and evaluate several instance sizes going from 100 to 1000. For the decoder layer depth scaling, we use CVRP100, 100 starting points and 160 sequential attempts. We then evaluate methods for a depth going from 1 to 10.
@@ Appendix B Training procedure
1:  Input: problem distribution $\mathcal{D}$, problem size $N$, memory $M$, batch size $B$, budget $K$, number of training steps $T$, policy $\mathbb{\pi}_{\theta}$ with pre-trained parameters $\theta$.
@@
10:        $m_{j}\leftarrow f(m_{j},\tau_{i}^{j})$   {Update the memory with transitions data}
@@
13:     $\nabla L(\tilde{\theta})\leftarrow\frac{1}{K}\sum_{i=1}^{K}\nabla L(\tilde{\theta})$   {Accumulate gradients}
14:     $\tilde{\theta}\leftarrow\tilde{\theta}-\alpha\nabla L(\tilde{\theta})$   {Update parameters}
@@ Appendix C Hyper-parameters
We report all the hyper-parameters used during train and inference time. For our method MEMENTO, there is no training hyper-parameters to report for instance sizes 125, 150, and 200 as the model used was trained on instances of size 100. The hyper-parameters used for MEMENTO are reported in Table 4. Since we also trained POMO and COMPASS on larger instances, we report hyper-parameters used for POMO and COMPASS in Table 5 and Table 6, respectively.
@@
num. CMAES components
CMAES init. sigma
@@ Appendix D Model checkpoints
Our experiments focus on two CO routing problems, TSP and CVRP, with methods being trained on two distinct instance sizes: 100 and 500. Whenever possible, we re-use existing checkpoints from the literature; in the remaining cases, we release all our newly trained checkpoints in the repository https://github.com/instadeepai/memento.
We evaluate MEMENTO on two CO problems, TSP and CVRP, and compare the performance to that of two main baselines: POMO (Kwon et al., 2020) and EAS (Hottung et al., 2022). The checkpoints used to evaluate POMO on TSP and CVRP are the same as the one used in Grinsztajn et al. (2023) and Chalumeau et al. (2023b), and the EAS baseline is executed using the same POMO checkpoint. These checkpoints were taken in the publicly available repository https://github.com/instadeepai/poppy. The POMO checkpoint is used in the initialisation step of MEMENTO (as described in Appendix B). To combine MEMENTO and COMPASS on TSP100, we re-use the COMPASS checkpoint available at https://github.com/instadeepai/compass and add the memory processing layers from the MEMENTO checkpoint trained on TSP100. This checkpoint is also released at https://github.com/instadeepai/memento.
For larger instances, we compare our MEMENTO method to three baselines: POMO, EAS and COMPASS. Since no checkpoint of POMO and COMPASS existed, we trained them with the tricks explained in Section 4. The process to generate the MEMENTO checkpoint and the MEMENTO (COMPASS) checkpoint is then exactly the same. All those checkpoints are available, for both TSP and CVRP.
@@ Appendix E Implementation details
The code-base in written in JAX (Bradbury et al., 2018), and is mostly compatible with recent repositories of neural solvers written in JAX, i.e. Poppy and COMPASS. The problems’ implementation are also written in JAX and fully jittable. Those come from the package Jumanji (Bonnet et al., 2023). CMA-ES implementation to mix MEMENTO and COMPASS is taken from the research package QDax (Chalumeau et al., 2023a). Neural networks, optimizers, and many utilities are implemented using the DeepMind JAX ecosystem (Babuschkin et al., 2020).
@@ Appendix F Can MEMENTO discover the REINFORCE update?
In Section 3.2, we presented the architecture used by the auxiliary model that processes the memory data to derive the new action logits. The intuition behind this architecture choice is that it should be able to learn the REINFORCE update rule. Indeed the REINFORCE loss associated with a new transition is $R\log(\pi_{\theta}(a))$, such that $\frac{\partial R\log(\pi_{\theta}(a))}{\partial l_{a}}=R(1-\pi_{\theta}(a))$. Therefore, simply having $H_{\theta_{M}}(f_{a})$ match $R(1-\pi_{\theta}(a))$ would recover a REINFORCE-like update. This is feasible, as $R$ and $\log(\pi_{\theta}(a))$ are included in the features $f_{a}$.
On Fig. 4, we compare the rule learned by MEMENTO compared to REINFORCE. In Appendix G, we present an ablation study of the features used in the update rule of MEMENTO, showing how much performance can be gained from the use of more information to derive the update rule.

### v3

@@ 5 Conclusion
We present memento, a method to improve adaptation of neural CO solvers to unseen instances by conditioning a policy directly on data collected online during search and search budget. In practice, it proves to outperform stochastic sampling, tree search and policy-gradient fine-tuning, and shows zero-shot combination with an unseen solver. Additionally, memento respects several key properties, like robustness to low-budget regimes, and favourable time performance scalability. We demonstrate the efficacy of memento by achieving SOTA single-policy adaptation on a standard benchmark on TSP and CVRP, both in and out-of-distribution. Moreover, we show that memento finds interpretable update rules to the underlying policy; trading off exploration and exploitation over the search budget and outperforming REINFORCE-style updates. We further demonstrate zero-shot combination of memento and compass, achieving overall SOTA on a benchmark of $12$ tasks.
Limitations and future work memento incurs additional compute and memory-usage compared to the memory-less base policies which it augments. Practically, we find that the performance gain significantly outweighs these overheads. A potential mitigation could be to derive the mathematical update learned by memento to avoid relying on the MLP computation. While we demonstrate successful zero-shot combination of memento with compass, training memento with a diversity-based method or fine-tuning memento specifically on the compass policy represents a promising direction for future work.
Acknowledgements
We thank the Python and JAX communities for developing tools that made this research possible. We thank the anonymous reviewers for their valuable comments and suggestions that helped improve the final version of this paper. Finally, we thank Google’s TPU Research Cloud (TRC) for supporting our research with Cloud TPUs.
@@ Appendix A Extended results
In Section 4, we compare memento to popular methods from the literature in numerous settings. In particular, Table˜1 compares using stochastic sampling, memento and eas to adapt pomo on the standard benchmark; and Figure 6 compares the relative performances of memento and eas on larger instances of TSP and CVRP for several values of batch sizes and budget. In this section, we report additional values and results of other methods for those experiments.
@@ A.1 Standard benchmark
We compare to the heuristic solver LKH3 [Helsgaun, 2017]; the current leading industrial solver of both TSP and CVRP, as well as an exact solver Concorde [Applegate et al., 2006] which is a TSP-specific industrial solver, and the CVRP-specific solver [Vidal et al., 2012].
We use datasets of 10,000 instances with 100 cities/customer nodes drawn from the training distribution, and three generalization datasets of 1,000 instances of sizes 125, 150, and 200, all from benchmark sets frequently used in the literature [Kool et al., 2019, Kwon et al., 2020, Hottung et al., 2022, Grinsztajn et al., 2023, Chalumeau et al., 2023b]. Table 5 displays results for TSP and CVRP on the standard benchmark. The columns provide the absolute tour length, the optimality gap, and the total inference time that each method takes to solve one instance within the attempts budget.
Standard benchmark with augmentation trick
Augmentation with symmetries is a problem-specific trick that can only be used for a few CO problems. Most prior work assumes that this additional x8 batching can be achieved seamlessly, which is unlikely in practice, when simulating complex real-world scenarios. Using this trick means decreasing the room for search and adaptation, since 87.5% of the budget is consumed to squeeze performance through uncontrolled network variance, rather than letting methods use principled strategies. Nevertheless, we report the standard benchmark with the "augmentation trick" in Table 6. It can be observed that memento outperforms eas in distribution (CVRP 100), and for the large instances task (CVRP 200). eas leads on the two other tasks. memento leads the augmented benchmark overall: it consistently leads in distribution, and both methods are competing closely out-of-distribution. It is to be noted that we have not tuned memento’s hyper-parameters for these runs.
@@
We were able to do so on the standard benchmark since we have implementations of most methods we were comparing in a single framework. Since we are reporting several external methods in Table˜8, we could only report time taken for the full dataset in that case.
Performance with runtime as budget
Expressing budget as time taken to solve one instance is subject to high biases but it is gives interesting perspectives on the time analysis and effects of parallelism. In Table 7, we provide results of using runtime in stead of attempts as budget given to each methods to solve TSP and CVRP instances. We use eas runtime as the reference time. These results may only be considered moderately since they are subject to a number of limitations: (i) different labs use different implementations and frameworks (ii) labs have access to different hardwares (iii) time is sensitive to parallelism, whereas number of attempts is not. Hence, comparing methods with time is subject to a higher number of biases, and would make it almost impossible to compare papers without a common codebase and hardware. We observe that memento still leads the benchmark, with 6 out of 8 top results; and both methods are very close on the two tasks where eas leads.
@@
On Table˜5, we can see that in the single-agent setting, memento leads the entire benchmark, and in the population-based setting, memento (compass) is leading the whole benchmark. memento is able to give significant improvement to compass on CVRP, pushing significantly the state-of-the-art on this benchmark. On TSP, the improvement is not significant enough to be visible on the rounded results. To finish with, we can observe that the time cost associated with memento is reasonable and is worth the performance improvement (except maybe for TSP results of memento (compass)).
@@
All neural methods reported in Table˜5 are our own runs with standardised checkpoints and rollout strategies, except for the results of sgbs, which were taken from Choo et al. [2022]. This introduces three biases: (i) the base pomo checkpoint used by sgbs is not exactly the same as our re-trained pomo checkpoint (ii) sgbs uses the domain-specific augmentation trick that we do not use (iii) sgbs pre-selects starting points in CVRP, which we do not do.
@@
Table 5: Results of memento against the baseline algorithms for (a) TSP and (b) CVRP. The methods are evaluated on instances from training distribution ($n=100$) as well as on larger instance sizes to test generalization. We use the same dataset as the rest of the literature, those contain 10 00010\,000 instances for $n=100$ and $1000$ instances for $n=125,150,200$. Gaps are computed relative to LKH3. We report time needed to solve one instance. sgbs * results are reported from Choo et al. [2022], and do not use the same pomo checkpoint as other reported results. Additionally, they rely on problem-specific tricks that were not used by other methods. Details in Section˜A.1.
@@
272S
Table 6: Results of memento and the baseline algorithms with instance augmentation for (a) TSP and (b) CVRP.
(a) TSP
Training distr. Generalization
$n=100$ $n=125$ $n=150$ $n=200$
Method Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time
sgbs 7.769 0.058% - - - - 9.367 0.220% - 10.753 0.619% -
pomo (sampling) 7.767 0.026% 16S 8.594 0.128% 20S 9.376 0.321% 29S 10.916 2.14% 45S
eas 7.768 0.038% 39S 8.590 0.080% 46S 9.361 0.159% 64S 10.730 0.403% 91S
memento (pomo) 7.765 0.008% 43S 8.586 0.035% 52S 9.355 0.091% 77S 10.743 0.526% 115S
(b) CVRP
Training distr. Generalization
$n=100$ $n=125$ $n=150$ $n=200$
Method Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time
HGS 15.563 $0.536\% -$ - - - 19.055 $0.884\% -$ 21.766 $1.096\% -$
LKH3 15.646 $0.000\%$ - 17.50 $0.000\%$ - 19.222 $0.000\%$ - 22.003 $0.000\%$ -
sgbs 15.659 0.08% - - - - 19.426 1.08% - 22.567 2.59% -
pomo (sampling) 15.67 0.18% 24S 17.56 0.33% 34S 19.43 1.08% 52S 23.24 5.64% 87S
eas 15.623 -0.175% 66S 17.473 -0.153% 82S 19.261 0.213% 123S 22.556 2.49% 179S
memento (pomo) 15.616 -0.196% 118S 17.511 0.040% 150S 19.316 0.477% 169S 22.515 2.308% 392S
Table 7: Results of memento and the baseline algorithms with budget expressed as runtime for (a) TSP and (b) CVRP.
(a) TSP
Training distr. Generalization
$n=100$ $n=125$ $n=150$ $n=200$
Method Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time
LKH3 7.765 0.000% - 8.583 0.000% - 9.346 0.000% - 10.687 0.000% -
pomo (sampling) 7.779 0.216% 39S 8.609 0.299% 46S 9.401 0.585% 64S 10.956 2.513% 91S
memento 7.768 0.046% 39S 8.592 0.110% 46S 9.365 0.203% 64S 10.760 0.681% 91S
(b) CVRP
Training distr. Generalization
$n=100$ $n=125$ $n=150$ $n=200$
Method Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time
LKH3 15.647 0.000% - 17.504 0.000% - 19.225 0.000% - 22.007 0.000% -
pomo (sampling) 15.713 0.399% 66S 17.612 0.642% 82S 19.488 1.393% 123S 23.378 6.264% 179S
memento 15.660 0.086% 66S 17.526 0.127% 82S 19.321 0.502% 123S 22.546 2.450% 179S
@@ A.2 Performance over different number of parallel and sequential attempts
In Section 4.3, we show performance improvement of memento over eas on instances of size 500 on TSP and CVRP for increasing number of sequential attempts and size of attempt batch. We report extended results with an additional competitor, pomo. We compare the online adaptation of the methods over four different sizes of batched solutions across increasing sequential attempts for each CO problem. Results from Table 8 show results of the three methods on TSP and CVRP. The columns show the best tour length performance for various values of sequential attempts expressed as budget, and solution batches of size $N$. Performance is averaged over a set of 128 instances.
@@ A.3 Evaluation over larger instances
In Section˜4.3, we evaluate memento and baselines on instances of size 500. For TSP, we use the dataset from Fu et al. [2021]. For CVRP, we use the dataset from Luo et al. [2023]. We do not include LEHD in our results since we focus on methods trained with Reinforcement Learning, and LEHD can only be successfully trained with supervised learning at the time of writing. Nevertheless, the good performance achieved by LEHD has motivated us to study the scaling law of memento as the number of layers in the decoder increases, reported on Fig.˜7.
@@
A.4 Time and memory complexity analysis
Time complexity
To get the time curves reported in Fig.˜7, we used CVRP, since it is the environment were memento was the slowest compared to eas. We hence expect curves on TSP to be even better for memento. For the instance size scaling, we use 50 starting points and 100 sequential attempts, and 1 layer in the decoder, and evaluate several instance sizes going from 100 to 1000. For the decoder layer depth scaling, we use CVRP100, 100 starting points and 160 sequential attempts. We then evaluate methods for a depth going from 1 to 10.
Computational and memory complexity analysis
The use of a memory to compute corrected action logits (i.e., memento’s policy update rule) introduces a computational overhead, whose scaling properties depend on specific design choices and problem parameters.
Consider the TSP100 example with the memory design presented in the paper. The memory contains 100 slots (one per node), each storing 40 entries with 5 float32 values, repeated for each starting point (100). This results in roughly 8 MB of memory data. The memory footprint therefore scales linearly with the instance size (since there is one slot per node), but remains independent of the base policy size. In addition, the MLP module used to process memory entries is small, consisting of two layers with eight hidden units.
Regarding computational cost, the main overhead comes from processing the retrieved entries through the MLP. At each step, 40 entries are retrieved and processed in parallel, so the effective cost is that of 40 independent MLP forward passes. This operation does not depend on the instance size or the policy’s parameter count, but only on the number of retrieved entries. An additional cost arises from indexing the appropriate slice of the memory, which can grow slightly with the number of nodes.
In practice, these overheads make memento slower than eas for small instance sizes or small neural networks. However, since its operations scale more favorably with instance size and parameter count (compared to backpropagation), memento becomes faster for larger problems.
Overall, the computational and memory overhead of memento is negligible when solving a few instances at a time, typical in practical settings where one receives a new problem instance and a limited compute budget, but it can become a bottleneck when solving large batches of problems in parallel, as often done in research benchmarks.
To provide a more concrete view of the memory-processing cost across instance sizes, we also report the runtime of pomo in Table˜9. The difference between the two indicates the additional cost introduced by memento. Note that these values may depend on the backend: for example, JAX’s jit compilation can impact relative timings, and alternative implementations (e.g., PyTorch) could yield different ratios.
Table 9: Average inference time (in seconds) for each algorithm across instance sizes.
Instance Size 100 250 500 750 875 950
pomo 57 114 225 360 443 504
eas 84 194 446 774 973 1122
memento 148 212 425 684 833 940
Figure 7: Time complexity of memento and eas for increasing values of decoder size, and instance size. memento is shown to scale better in time than eas.
@@ Appendix B Training procedure
1:Input: problem distribution $\mathcal{D}$, problem size NN, memory $M$, batch size $B$, budget $K$, number of training steps $T$, policy $\mathbb{\pi}_{\theta}$ with pre-trained parameters $\theta$.
@@
10:   $m_{j}\leftarrow f(m_{j},\tau_{i}^{j})$   {Update the memory with transition data}
@@
[…]
@@ Appendix C Hyper-parameters
We report all the hyper-parameters used during train and inference time. For our method memento, there is no training hyper-parameters to report for instance sizes 125, 150, and 200 as the model used was trained on instances of size 100. The hyper-parameters used for memento are reported in Table˜10. Since we also trained pomo and compass on larger instances, we report hyper-parameters used for pomo and compass in Table˜11 and Table˜12, respectively.
@@
num. cma-es components
cma-es init. sigma
@@ Appendix D Model checkpoints
Our experiments focus on two CO routing problems, TSP and CVRP, with methods being trained on two distinct instance sizes: 100 and 500. Whenever possible, we re-use existing checkpoints from the literature; in the remaining cases, we release all our newly trained checkpoints in the repository anonymised for the review process.
We evaluate memento on two CO problems, TSP and CVRP, and compare the performance to that of two main baselines: pomo [Kwon et al., 2020] and eas [Hottung et al., 2022]. The checkpoints used to evaluate pomo on TSP and CVRP are the same as the one used in Grinsztajn et al. [2023] and Chalumeau et al. [2023b], and the eas baseline is executed using the same pomo checkpoint. These checkpoints were taken in the publicly available repository https://github.com/instadeepai/poppy. The pomo checkpoint is used in the initialisation step of memento (as described in Appendix˜B). To combine memento and compass on TSP100, we re-use the compass checkpoint available at https://github.com/instadeepai/compass and add the memory processing layers from the memento checkpoint trained on TSP100. This checkpoint is also released at anonymised for the review process.
For larger instances, we compare our memento method to three baselines: pomo, eas and compass. Since no checkpoint of pomo and compass existed, we trained them with the tricks explained in Section˜4. The process to generate the memento checkpoint and the memento (compass) checkpoint is then exactly the same. All those checkpoints are available, for both TSP and CVRP.
@@ Appendix E Implementation details
The code-base in written in JAX [Bradbury et al., 2018], and is mostly compatible with recent repositories of neural solvers written in JAX, i.e. Poppy and compass. The problems’ implementation are also written in JAX and fully jittable. Those come from the package Jumanji [Bonnet et al., 2023]. CMA-ES implementation to mix memento and compass is taken from the research package QDax [Chalumeau et al., 2023a]. Neural networks, optimizers, and many utilities are implemented using the DeepMind JAX ecosystem [Babuschkin et al., 2020].
@@ Appendix F Can MEMENTO discover the REINFORCE update?
In Section˜3.2, we presented the architecture used by the auxiliary model that processes the memory data to derive the new action logits. The intuition behind this architecture choice is that it should be able to learn the REINFORCE update rule. Indeed the REINFORCE loss associated with a new transition is $R\log(\pi_{\theta}(a))$, such that $\frac{\partial R\log(\pi_{\theta}(a))}{\partial l_{a}}=R(1-\pi_{\theta}(a))$. Therefore, simply having $H_{\theta_{M}}(f_{a})$ match $R(1-\pi_{\theta}(a))$ would recover a REINFORCE-like update. This is feasible, as $R$ and $\log(\pi_{\theta}(a))$ are included in the features $f_{a}$.
We compare the rule learned by memento to REINFORCE in Fig.˜4 and provide an analysis of how the rule learned by memento evolve over different budgets in Appendix˜J. In Appendix˜G, we present an ablation study of the features used in the update rule of memento, showing how much performance can be gained from the use of more information to derive the update rule.